I need to make this equation as an check answer
18x+2(x-100=3500
18x+2x-200=3500
18x 2x=3700
18x=185
This is the problem but I don't know who to check the problem

Answer by oscargut(2103)

18x+2(x-100)=3500
18x+2x-200=3500
20x = 3700
x = 185

Checking:
18(185)+2(185-100) = 3500
3330+2(85) = 3500
3330+170 = 3500
3500 = 3500
I can help you with more questions at: mthman@gmail.com
Thanks